Session: May 2003 (session 2)
Wavelength: 5 cm
To give feedback on this experiment, click
Medicina (A. Cattani at Fri Jun 6 21:13:51 2003) Success :
Hartebeesthoek (Jonathan Quick at Fri Jun 6 21:27:03 2003) Success :
Onsala (Michael Lindqvist at Sat Jun 7 06:55:45 2003) Minor failures :
Could only observe at 6500 MHz during an experiment.
Torun (K.Borkowski at Mon Jun 9 05:16:47 2003) Success :
No known problems.
Effelsberg (A. Kraus at Tue Jun 10 09:14:59 2003) Minor failures : RECO
Tape drive problems in the beginning led to the loss of data between UT1700-1717 and UT1732-1749.
Jodrell_Bank (Alastair Gunn at Tue Jun 10 12:58:36 2003) Success :
No known problems. Changed procedure file to reset LOs as required.
Cambridge (Alastair Gunn at Tue Jun 10 12:59:17 2003) Success :
No known problems. Changed procedure file to set LOs as required.